# Diffusion Models for Image Restoration and Enhancement: A Comprehensive Survey

### v2

@@ Abstract
Keywords:
Diffusion models Image restoration Image enhancement Image processing
@@ 1 Introduction
Image Restoration (IR) has been a long-term research topic in low-level vision tasks, which plays an irreplaceable role in improving the subjective quality of images. The popular IR tasks contain image super-resolution (SR) (Dong et al., 2014; Lim et al., 2017a; Dai et al., 2019; Wang et al., 2018a; Lu et al., 2022; Chen et al., 2023b; Zhang et al., 2022a; Gao et al., 2023a), deblurring (Nah et al., 2017a; Kupyn et al., 2018; Nimisha et al., 2017; Kim et al., 2023b; Zamir et al., 2022; Tsai et al., 2022), denoising (Zhang et al., 2017; Valsesia et al., 2020; Fan et al., 2022; Liu et al., 2022a; Tian et al., 2023; Zhang et al., 2023a), inpainting (Xie et al., 2012; Pathak et al., 2016; Zhao et al., 2020; Yu et al., 2023b; Li et al., 2022c) and compression artifacts removing (Dong et al., 2015; Svoboda et al., 2016; Ehrlich et al., 2020; Wang et al., 2022b; Jiang et al., 2023c), etc. To restore the distorted images, traditional IR methods treated the restoration as the signal processing and reduce the artifacts with hand-crafted algorithms from the spatial or frequency perspective (Dabov et al., 2007; Chantas et al., 2009; Mairal et al., 2008; Xu and Sun, 2010; Schmidt and Roth, 2014). With the development of deep learning, numerous IR works collected a series of datasets tailored for various IR tasks, e.g., DIV2K (Agustsson and Timofte, 2017), LSDIR (Li et al., 2023d), Set5 (Bevilacqua et al., 2012), and Set14 (Zeyde et al., 2012) for SR, Rain800 (Zhang et al., 2019a), Rain200 (Yang et al., 2017), Raindrop (Qian et al., 2018) and DID-MDN (Zhang and Patel, 2018) for draining, REDS (Nah et al., 2019), and Gopro (Nah et al., 2017b) for motion deblurring, etc. Leveraging these datasets, the majority of recent works (Lim et al., 2017a; Nah et al., 2017a; Valsesia et al., 2020; Dong et al., 2015; Svoboda et al., 2016; Zhang et al., 2017; Lu et al., 2022; Liu et al., 2021; Zhang et al., 2022a; Wang et al., 2022e; Zamir et al., 2022) focused on improving the representation capability of IR networks for complicated degradation through well-designed backbones based on Convolutional neural networks (CNNs) (Krizhevsky et al., 2017) or Transformer (Vaswani et al., 2017). Although these works achieve great progress in the objective quality (e.g., PSNR, and SSIM), the restored images still suffer from unsatisfactory texture generation, hindering the application of these IR methods in real-world scenarios.
@@
Thanks to the development of generative models (Kingma and Welling, 2013; Van Den Oord et al., 2016; Salimans et al., 2017; Rezende and Mohamed, 2015; Papamakarios et al., 2017; Creswell et al., 2018; Karras et al., 2017a), especially the generative adversarial network (GAN) (Creswell et al., 2018), some pioneering IR studies (Dosovitskiy and Brox, 2016; Yu and Porikli, 2016; Johnson et al., 2016; Ledig et al., 2017; Wang et al., 2018a) pinpoint that previous pixel-wise loss, e.g., MSE loss, and L1 loss, are susceptible to the blurring textures, and introduce the adversarial loss from GAN to the optimization of IR network, for enhancing its texture generation capability. For instance, SRGAN (Ledig et al., 2017) and DeblurGAN (Kupyn et al., 2018) utilize a combination of pixel-wise loss and adversarial loss to achieve perception-oriented SR network and deblurring network, respectively. Following them, two primary directions to improve GAN-based IR arise by enhancing the generator (i.e., the restoration network) (Ledig et al., 2017; Zhang et al., 2019b; Wang et al., 2018a, 2021b; Zhang et al., 2021a) and discriminator (Schonfeld et al., 2020; Isola et al., 2017; Karras et al., 2020). In particular, ESRGAN (Wang et al., 2018a) introduces the powerful RRDB (Wang et al., 2018a) as a generator for GAN-based SR tasks. Three popular discriminators, including pixel-wise discriminator (U-Net shape) (Schonfeld et al., 2020), patch-wise discriminator (Isola et al., 2017; Wang et al., 2018b), and image-wise discriminator (Simonyan and Zisserman, 2014; Karras et al., 2020) (i.e., VGG-like architecture) are designed to focus on the subjective quality at different levels of granularity, (i.e., from local to global). Although the above progress, most researches on GAN-based IR still face two inevitable but crucial problems: 1) the training of GAN-based IR is susceptible to mode corruption and unstable optimization and 2) the textures of most generated images seem to be fake and counterfactual.
@@
Recently, diffusion models have emerged as a new branch of generative models, leading to a series of breakthroughs in visual generation tasks. The prototype of the diffusion model can be traced back to the work (Sohl-Dickstein et al., 2015a), and has been developed by DDPM (Ho et al., 2020), NCSN (Song and Ermon, 2019) and SDE (Song et al., 2020b). In general, the diffusion model is composed of the forward/diffusion process and reverse process, where the forward process progressively increases the pixel-wise noise to the image until it satisfies the distribution of Gaussian noise, and the reverse process aims to reconstruct the image by denoising with score estimating (Song and Ermon, 2019) or noise prediction (Ho et al., 2020). Compared to GANs, the diffusion models yield high-fidelity and diverse generation results, thereby successfully replacing the GANs in a series of fields, such as visual generation (Song and Ermon, 2019; Song et al., 2020a, b; Ho et al., 2020) and conditional visual generation (Singh et al., 2022; Popov et al., 2021; Babnik et al., 2023; Blattmann et al., 2023). With the advancement of the vision-language model, diffusion model has been extended to the cross-modality generation, such as StableDiffusion (Rombach et al., 2022), and DALLE-2 (Ramesh et al., 2022). This greatly promotes the development of artificial intelligence generated content (AIGC). We have listed the representative works on diffusion models based on the timeline in Fig. 1 of the Supplementary.
@@
Inspired by the superior generative capability of diffusion models, numerous studies have investigated their application in image restoration tasks, targeting to facilitate the texture recovery. According to the training strategy, these works can be roughly divided into two categories: 1) the first category (Saharia et al., 2022b; Li et al., 2022b; Dos Santos et al., 2022; Özdenizci and Legenstein, 2023; Whang et al., 2022; Luo et al., 2023c; Zhou et al., 2023; Chan et al., 2023; Miao et al., 2023; Varanka et al., 2024; Qu et al., 2024; Zhang et al., 2023d) is dedicated to optimizing the diffusion model for IR from scratch via supervised learning, and 2) the second one (i.e., zero-shot one) (Kawar et al., 2021, 2022b, 2022a; Wang et al., 2023g; Chung et al., 2023b; Song et al., 2021; Liu et al., 2023b; Abu-Hussein et al., 2022) strives to exploit the generative priors in the pre-trained diffusion models for IR. Typically, supervised learning-based methods necessitate the collection of large-scale distorted/clean image pairs, while zero-shot-based methods predominantly rely on known degradation modes. These limitations impede the application of these diffusion model-based methods in real-world scenarios, where the distortions are typically diverse and unknown. To further tackle the above issue, some researches (Sahak et al., 2023; Murata et al., 2023; Chung et al., 2023a; Wang et al., 2023c; Wei et al., 2023; Wang et al., 2023i) have extended the diffusion model to handle blind/real-world image restoration by incorporating real-world distortion simulation, kernel estimation, domain translation, and distortion invariant learning.
@@
In this survey, we start by presenting the background of diffusion models in Sec. 2, highlighting two foundational modeling methods, i.e., DDPM (Ho et al., 2020), and SDE (Song et al., 2020b). Based on these preliminaries, we shed light on the advances of diffusion models in image restoration from two distinct directions in Sec. 3: 1) supervised diffusion model-based IR, and 2) zero-shot diffusion model-based IR. In Sec. 4, we summarize the diffusion model-based IR under more practical and challenging scenarios, i.e., blind/real-world degradation. This intends to further enhance the capability of the diffusion model-based IR methods in fulfilling the demands of practical application. To facilitate a reasonable and exhaustive comparison, in Sec. 5, a series of comparisons between different DM-based benchmarks across four tasks are provided, including image SR, inpainting, and deblurring, and real-world restoration. In Sec. 6, we delve into analyzing the primary challenges and potential directions in diffusion model-based IR to provide some insights and hope to inspire more new works. The final conclusion for this review is summarized in Sec. 7.
@@ 2 Background on Diffusion Model (DM)
Diffusion probabilistic model (i.e., diffusion model) has brought an evolution in the field of generative models, which transforms the complicated and unstable generation process into several independent and stable reverse processes via Markov Chain modeling. There are three foundational diffusion models that are widely utilized, including DDPM (Ho et al., 2020). NCSNs (Song and Ermon, 2019) and SDE (Song et al., 2020b). Among them, NCSNs (Song and Ermon, 2019) seeks to model the data distribution by sampling using annealed Lange-vin dynamics with a sequence of decreasing noise scales. In contrast, DDPM (Ho et al., 2020) models the forward process with a fixed process of adding Gaussian noise, which simplifies the reverse process of the diffusion model into a solution process for the variational bound objective. These two basic diffusion models are actually special cases of score-based generative models (Song et al., 2020b). SDE (Song et al., 2020b), as the unified form, models the continuous diffusion and reverse processes with stochastic differential equation (SDE). It proves that the NCSNs and DDPM are only two separate discretizations of SDE. We will clarify two commonly-used modeling strategies of diffusion models in image restoration, i.e., DDPM (Ho et al., 2020) and SDE (Song et al., 2020b) in the following subsections. We also illustrate the NCSN (Song and Ermon, 2019) in the section B.1 of the Supplementary.
@@ 2.1 Denoising Diffusion Probabilistic Model
DDPM (Denosing diffusion probabilistic model) (Ho et al., 2020) originates from the diffusion models (Sohl-Dickstein et al., 2015b), which introduces the simple variational bound objective for diffusion models by setting the variance $\beta_{t}$ as fixed values. There are two crucial processes in diffusion models, i.e., the forward process and reverse process. In particular, the forward process (i.e., the diffusion process in DDPM) aims to progressively corrupt the training data to the Gaussian Noise, which is a parameterized Markov chain as:
@@
where $\tilde{\mu}_{t}(x_{t},x_{0})=\frac{\sqrt{\hat{\alpha}_{t-1}}\beta_{t}}{1-\hat{\alpha}_{t}}x_{0}+\frac{\sqrt{\hat{\alpha}_{t}}(1-\hat{\alpha}_{t-1})}{1-\hat{\alpha}_{t}}x_{t}=\frac{1}{\sqrt{\alpha_{t}}}(x_{t}\\ -\frac{\beta_{t}}{\sqrt{1-\hat{\alpha}_{t}}})\epsilon$ for $\epsilon\sim\mathcal{N}(\mathbf{0},\textbf{I})$ and $\tilde{\beta_{t}}=\frac{1-\hat{\alpha}_{t-1}}{1-\hat{\alpha}_{t}}$. As stated in Eq. 3, the variance schedule $\beta_{t}$ is predefined, and thus, it only requires approximating the mean $\mu_{\theta}(x_{t},t)=\tilde{\mu}_{t}(x_{t},x_{o})$ by a denoising network $\epsilon_{\theta}(x_{t},t)$. The optimization objective (Ho et al., 2020) for denoising network can be written as:
@@ 2.2 Stochastic Differential Equations (SDEs)
To unify the approaches with the score-based generative modeling and diffusion probabilistic modeling, SDEs (Song et al., 2020b) exploit the continuous diffusion process through the stochastic differentiable equation (SDE) as:
@@
where $\mathbf{w}$ is the standard Wiener process, $\mathbf{f}(\cdot,t)$ is called drift coefficient of $x(t)$ and $g(\cdot)$ is called the diffusion coefficient of $x(t)$. Here, the diffusion coefficient can be understood as the degree perturbed by random noise, and the drift coefficients can be designed to ensure the Gaussian distribution, such as DDPM (Ho et al., 2020) and NCSN (Song and Ermon, 2019). The reverse process of the above continuous diffusion process (i.e., sampling data from noise) is also a diffusion process and can be modeled by a reverse-time SDE:
@@
Here $dt$ is an infinitesimal negative time step and $\hat{\mathbf{w}}$ is the standard Wiener process when time flows backward from $T$ to 0. The core of reverse-time SDE is to estimate the score function using a neural network and then Eq. 6 can be solved with the score matching (Hyvärinen, 2007; Hyvärinen and Dayan, 2005).

### v3

@@ 1 Introduction
Image Restoration (IR) has been a long-term research topic in low-level vision tasks, which plays an irreplaceable role in improving the subjective quality of images. The popular IR tasks contain image super-resolution (SR) (SR1_SRCNN-cnn, ; SR3_EDSR-cnn, ; SR4_SAN, ; wang2018esrgan, ; sr_transformer1, ; chen2023activating-HST, ; sr3_swinfir-transformer, ; gao2023ctcnet, ), deblurring (nah2017deep-deblur-old1-cnn, ; kupyn2018deblurgan-old2, ; nimisha2017blur-old3-cnn, ; kim2023mssnet-deblur-cnn, ; zamir2022restormer_deblur-transformer, ; tsai2022stripformer_deblur-transformer, ), denoising (zhang2017beyond-denoise2-cnn, ; valsesia2020deep-denoise4-cnn, ; fan2022sunet-denoise-new1-transformer, ; liu2022dnt-denoise-new2-transformer, ; tian2023multi-denoise-new3, ; zhang2023mm-new4, ), inpainting (xie2012image-inpainting-old1, ; pathak2016context-inpainting-old2, ; inpainting-new2, ; yu2023inpaint-new3, ; li2022mat-inpainting-new4, ) and compression artifacts removing (dong2015compression-jpeg-old1-cnn, ; svoboda2016compression-jpeg-old2-cnn, ; ehrlich2020quantization-jpeg-old3-cnn, ; wang2022jpeg-new2, ; jiang2023multi-jpeg-new3, ), etc. To restore the distorted images, traditional IR methods treated the restoration as the signal processing and reduce the artifacts with hand-crafted algorithms from the spatial or frequency perspective (dabov2007image-traditional1, ; traditional3, ; mairal2008learning-traditional-4, ; xu2010image-traditional-5, ; schmidt2014shrinkage-traditional-8, ). With the development of deep learning, numerous IR works collected a series of datasets tailored for various IR tasks, e.g., DIV2K (DIV2K, ), LSDIR (li2023lsdir, ), Set5 (set5, ), and Set14 (Set14, ) for SR, Rain800 (Rain800, ), Rain200 (yang2017deep-rain100, ), Raindrop (rain-drop, ) and DID-MDN (zhang2018density-DID-MDN, ) for draining, REDS (Nah_2019_CVPR_Workshops_REDS, ), and Gopro (Gopro, ) for motion deblurring, etc. Leveraging these datasets, the majority of recent works (SR3_EDSR-cnn, ; nah2017deep-deblur-old1-cnn, ; valsesia2020deep-denoise4-cnn, ; dong2015compression-jpeg-old1-cnn, ; svoboda2016compression-jpeg-old2-cnn, ; zhang2017beyond-denoise2-cnn, ; sr_transformer1, ; swinIR-transformer, ; sr3_swinfir-transformer, ; wang2022uformer-transformer, ; zamir2022restormer_deblur-transformer, ) focused on improving the representation capability of IR networks for complicated degradation through well-designed backbones based on Convolutional neural networks (CNNs) (krizhevsky2017imagenet-cnn1, ) or Transformer (vaswani2017attention-transformer, ). Although these works achieve great progress in the objective quality (e.g., PSNR, and SSIM), the restored images still suffer from unsatisfactory texture generation, hindering the application of these IR methods in real-world scenarios.
@@
Thanks to the development of generative models (kingma2013auto-VAE, ; van2016pixel-autoregressive1, ; salimans2017pixelcnn++-autoregressive3, ; rezende2015variational-flow, ; papamakarios2017masked-flow2, ; GAN, ; karras2017progressive-GAN1, ), especially the generative adversarial network (GAN) (GAN, ), some pioneering IR studies (dosovitskiy2016generating-loss1, ; yu2016ultra-loss2, ; johnson2016perceptual-loss3, ; SRGAN, ; wang2018esrgan, ) pinpoint that previous pixel-wise loss, e.g., MSE loss, and L1 loss, are susceptible to the blurring textures, and introduce the adversarial loss from GAN to the optimization of IR network, for enhancing its texture generation capability. For instance, SRGAN (SRGAN, ) and DeblurGAN (kupyn2018deblurgan-old2, ) utilize a combination of pixel-wise loss and adversarial loss to achieve perception-oriented SR network and deblurring network, respectively. Following them, two primary directions to improve GAN-based IR arise by enhancing the generator (i.e., the restoration network) (SRGAN, ; zhang2019ranksrgan-gene1, ; wang2018esrgan, ; wang2021realesrgan, ; zhang2021designing-gene2, ) and discriminator (schonfeld2020u-pixel-discri, ; isola2017image-patch-wise-discri, ; karras2020analyzing-image-wise-discri2, ). In particular, ESRGAN (wang2018esrgan, ) introduces the powerful RRDB (wang2018esrgan, ) as a generator for GAN-based SR tasks. Three popular discriminators, including pixel-wise discriminator (U-Net shape) (schonfeld2020u-pixel-discri, ), patch-wise discriminator (isola2017image-patch-wise-discri, ; wang2018fully-patch-wise-4, ), and image-wise discriminator (simonyan2014very-image-wise-discri, ; karras2020analyzing-image-wise-discri2, ) (i.e., VGG-like architecture) are designed to focus on the subjective quality at different levels of granularity, (i.e., from local to global). Although the above progress, most researches on GAN-based IR still face two inevitable but crucial problems: 1) the training of GAN-based IR is susceptible to mode corruption and unstable optimization and 2) the textures of most generated images seem to be fake and counterfactual.
@@
Recently, diffusion models have emerged as a new branch of generative models, leading to a series of breakthroughs in visual generation tasks. The prototype of the diffusion model can be traced back to the work (sohl2015deep-dpm, ), and has been developed by DDPM (DDPM, ), NCSN (NCSN, ) and SDE (SDE, ). In general, the diffusion model is composed of the forward/diffusion process and reverse process, where the forward process progressively increases the pixel-wise noise to the image until it satisfies the distribution of Gaussian noise, and the reverse process aims to reconstruct the image by denoising with score estimating (NCSN, ) or noise prediction (DDPM, ). Compared to GANs, the diffusion models yield high-fidelity and diverse generation results, thereby successfully replacing the GANs in a series of fields, such as visual generation (NCSN, ; DDIM, ; SDE, ; DDPM, ) and conditional visual generation (singh2022conditioning-noisecondition, ; popov2021grad-app1, ; babnik2023diffiqa-app11, ; blattmann2023align-app12, ). With the advancement of the vision-language model, diffusion model has been extended to the cross-modality generation, such as StableDiffusion (LDM, ), and DALLE-2 (ramesh2022hierarchical-dalle2, ). This greatly promotes the development of artificial intelligence generated content (AIGC). We have listed the representative works on diffusion models based on the timeline in Fig. 1 of the Supplementary.
@@
Inspired by the superior generative capability of diffusion models, numerous studies have investigated their application in image restoration tasks, targeting to facilitate the texture recovery. According to the training strategy, these works can be roughly divided into two categories: 1) the first category (SR3, ; li2022srdiff, ; sde-sr, ; weather-diff, ; deblur-DPM, ; luo2023refusion, ; zhou2023pyramid-pydiff, ; chan2023sud, ; miao2023dds2m, ; varanka2024pfstorer, ; qu2024xpsr, ; zhang2023diffusion-difftsr, ) is dedicated to optimizing the diffusion model for IR from scratch via supervised learning, and 2) the second one (i.e., zero-shot one) (kawar2021snips, ; jpegddrm, ; DDRMDRM, ; ddnm, ; dps, ; SIM-SGM, ; liu2023improved, ; abu2022adir, ) strives to exploit the generative priors in the pre-trained diffusion models for IR. Typically, supervised learning-based methods necessitate the collection of large-scale distorted/clean image pairs, while zero-shot-based methods predominantly rely on known degradation modes. These limitations impede the application of these diffusion model-based methods in real-world scenarios, where the distortions are typically diverse and unknown. To further tackle the above issue, some researches (SR3+, ; murata2023gibbsddrm, ; blinddps, ; wang2023exploiting-stable-sr, ; wei2023raindiffusion, ; wang2023dr2, ) have extended the diffusion model to handle blind/real-world image restoration by incorporating real-world distortion simulation, kernel estimation, domain translation, and distortion invariant learning.
@@
In this survey, we start by presenting the background of diffusion models in Sec. 2, highlighting two foundational modeling methods, i.e., DDPM (DDPM, ), and SDE (SDE, ). Based on these preliminaries, we shed light on the advances of diffusion models in image restoration from two distinct directions in Sec. 3: 1) supervised diffusion model-based IR, and 2) zero-shot diffusion model-based IR. In Sec. 4, we summarize the diffusion model-based IR under more practical and challenging scenarios, i.e., blind/real-world degradation. This intends to further enhance the capability of the diffusion model-based IR methods in fulfilling the demands of practical application. To facilitate a reasonable and exhaustive comparison, in Sec. 5, a series of comparisons between different DM-based benchmarks across four tasks are provided, including image SR, inpainting, and deblurring, and real-world restoration. In Sec. 6, we delve into analyzing the primary challenges and potential directions in diffusion model-based IR to provide some insights and hope to inspire more new works. The final conclusion for this review is summarized in Sec. 7.
@@ 2 Background on Diffusion Model (DM)
Diffusion probabilistic model (i.e., diffusion model) has brought an evolution in the field of generative models, which transforms the complicated and unstable generation process into several independent and stable reverse processes via Markov Chain modeling. There are three foundational diffusion models that are widely utilized, including DDPM (DDPM, ). NCSNs (NCSN, ) and SDE (SDE, ). Among them, NCSNs (NCSN, ) seeks to model the data distribution by sampling using annealed Lange-vin dynamics with a sequence of decreasing noise scales. In contrast, DDPM (DDPM, ) models the forward process with a fixed process of adding Gaussian noise, which simplifies the reverse process of the diffusion model into a solution process for the variational bound objective. These two basic diffusion models are actually special cases of score-based generative models (SDE, ). SDE (SDE, ), as the unified form, models the continuous diffusion and reverse processes with stochastic differential equation (SDE). It proves that the NCSNs and DDPM are only two separate discretizations of SDE. We will clarify two commonly-used modeling strategies of diffusion models in image restoration, i.e., DDPM (DDPM, ) and SDE (SDE, ) in the following subsections. We also illustrate the NCSN (NCSN, ) in the section B.1 of the Supplementary.
@@ 2.1 Denoising Diffusion Probabilistic Model
DDPM (Denosing diffusion probabilistic model) (DDPM, ) originates from the diffusion models (sohl2015deepDiffusionModels, ), which introduces the simple variational bound objective for diffusion models by setting the variance $\beta_{t}$ as fixed values. There are two crucial processes in diffusion models, i.e., the forward process and reverse process. In particular, the forward process (i.e., the diffusion process in DDPM) aims to progressively corrupt the training data to the Gaussian Noise, which is a parameterized Markov chain as:
@@
where $\tilde{\mu}_{t}(x_{t},x_{0})=\frac{\sqrt{\hat{\alpha}_{t-1}}\beta_{t}}{1-\hat{\alpha}_{t}}x_{0}+\frac{\sqrt{\hat{\alpha}_{t}}(1-\hat{\alpha}_{t-1})}{1-\hat{\alpha}_{t}}x_{t}=\frac{1}{\sqrt{\alpha_{t}}}(x_{t}\\ -\frac{\beta_{t}}{\sqrt{1-\hat{\alpha}_{t}}})\epsilon$ for $\epsilon\sim\mathcal{N}(\mathbf{0},\textbf{I})$ and $\tilde{\beta_{t}}=\frac{1-\hat{\alpha}_{t-1}}{1-\hat{\alpha}_{t}}$. As stated in Eq. 3, the variance schedule $\beta_{t}$ is predefined, and thus, it only requires approximating the mean $\mu_{\theta}(x_{t},t)=\tilde{\mu}_{t}(x_{t},x_{o})$ by a denoising network $\epsilon_{\theta}(x_{t},t)$. The optimization objective (DDPM, ) for denoising network can be written as:
@@ 2.2 Stochastic Differential Equations (SDEs)
To unify the approaches with the score-based generative modeling and diffusion probabilistic modeling, SDEs (SDE, ) exploit the continuous diffusion process through the stochastic differentiable equation (SDE) as:
@@
where $\mathbf{w}$ is the standard Wiener process, $\mathbf{f}(\cdot,t)$ is called drift coefficient of $x(t)$ and $g(\cdot)$ is called the diffusion coefficient of $x(t)$. Here, the diffusion coefficient can be understood as the degree perturbed by random noise, and the drift coefficients can be designed to ensure the Gaussian distribution, such as DDPM (DDPM, ) and NCSN (NCSN, ). The reverse process of the above continuous diffusion process (i.e., sampling data from noise) is also a diffusion process and can be modeled by a reverse-time SDE:
@@
Here $dt$ is an infinitesimal negative time step and $\hat{\mathbf{w}}$ is the standard Wiener process when time flows backward from $T$ to 0. The core of reverse-time SDE is to estimate the score function using a neural network and then Eq. 6 can be solved with the score matching (hyvarinen2007some-score-matching, ; hyvarinen2005estimation-score-matching, ).